OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. tn = 1/2, 1 3/4, 3, 4 1/4, ... tn = 5/4 + 1/2(n – 1) tn = 1/2 – 5/4(n – 1) tn = 1/2 + 5/4(n – 1) tn = 1/2 + 3(n – 5/4)

OpenStudy (anonymous):

See, General Formula for AP is given As: \[t_n = a + (n - 1).d\] Here, tn = Last term of AP, n = Number of terms in AP a = First term of AP, d = Common Difference of AP..

OpenStudy (cwrw238):

common difference = 1 3/4 - 1/2 = 5/4

OpenStudy (anonymous):

tn = a+ (n-1)d so 1/2+5/4(n-1)

OpenStudy (anonymous):

the series can be rewritten as 2/4, 7/4, 12/4, 17/4... so d= 5/4.. a= first term=1/2

OpenStudy (anonymous):

So, the terms given are: 1/2, 7/4, 17/4, ..... d= 7/4 - 1/2 = 5/4 Or you can further check it: 3 - 7/4 = 5/4 So, the d = 5/4 and a = 1/2 Put these values in the formula written above and solve for tn...

OpenStudy (anonymous):

Here I missed 3 in the terms given are...